STANDARD QUALITY COFFEE FOR $18.00/KG AND PRIME QUALITY COFE SELLS FOR $24.00/KG. WHAT QUANTITY OF EACH SHOULD BE COMBINED TO PRODUCE 40KG OF A BLEND TO SELL FOR $22.50/KG?
Answer by checkley71(8403)

24x+18(40-x)=22.50*40
24x+720-18x=900
6x=900-720
6x=180
x=180/6
x=30 kg of $24 coffee is needed.
40-30=10 kg of $18 coffee is needed.
proof
24*30=18*10=900
720+180=900
900=900